# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v3

@@ 3.2 Quantum Supremacy for Solving Complex Problems
The inherent problems’ difficulty, or hardness, is a fundamental concept in computational complexity theory (see Appendix A for more details). The hardness quantifies the resources required by different models of computation to solve problems, such as the number of steps of a deterministic Turing machine (Papadimitriou & Yannakakis, 1991). Various models and resources are often considered, including deterministic, nondeterministic, and probabilistic models, time and space constraints, and interactions across different models. “Quantum Supremacy” is a concept that indicates that a programmable quantum device can solve a problem that no classical computer can solve in a feasible amount of time. For instance, Google’s Sycamore processor (Frank Arute, 2019) performs in $200$ seconds a task that would require $10,000$ years using a classical computer. Such an acclaimed advantage of QC compared to classical computers demands a modification of the complexity theory for problems that are solvable on a quantum computer. Hence, the classical complexity classification diagram has been redefined to incorporate the quantum perspective and evaluate the complexity of the issues, considering their ability to be solvable on a quantum computer. Figure 13 shows that a set of problems outside the P space that can be solved in Bounded-error, Quantum, and Polynomial-time (BQP) (Watrous, 2009). Identifying BQP problems is crucial to prove the overwhelming ability of quantum computers compared to classical computers.
3.3 Quantum Enhancement to Classical Problems
In the NISQ era, as important as it is to lead the direction to showcase QC’s ability to outperform classical computers to solve problems with unattainable solutions within the classical paradigm. It is also necessary to pave a direction where we put both paradigms together to complement each other. It is a key step to comprehend how to use quantum and classical paradigms in an amalgamated approach to solutions that utilize the strengths of each framework in their own aspects and practicality. In contrast to the criteria set for achieving quantum advantage, the quantum enhancement’s goal is to show QC solving a hard problem in a realistic time frame, irrespective of the reliability of the solution. It is also important to present a stringent version of the passing criteria that dictate and drive the practical usefulness of Quantum Computers. Frameworks and solutions designed under such a lens can be seen as quantum enhancements to existing but not limited to current classical systems and designs.
Demonstrating such aspects of QC is essential to gain confidence and value in using such systems. Hence, works like Pokharel & Lidar (2023) that aim to practically demonstrate enhancements like quantum speedup for a computational problem are good examples of paths leading to practical quantum usability. Similar works investigate how the generic quantum advantage passing criteria can be tuned slightly towards strict reasoning to generate useful results and efforts.
@@ 3.3 Quantum Enhancement to Classical Problems
To be critical of the current state of quantum practicality, existing methods are either heuristic-based or depend largely on the encoding of data in a quantum mechanical way that causes the change in performance and are the sole aspects making the difference from classical performance. In light of the hybrid quantum-classical systems, the work of Liu et al. (2021) discusses important aspects that should be changed to obtain meaningful developments and revolves around how heuristic-based algorithms do not provide formal evidence that showcases their genuine and consistent advantage over classical algorithms. The work of Kashif et al. (2024) demonstrates that HQNNs exhibit better scalability than classical neural networks when increasing classification task complexity. Moreover, the variational circuits can only implement linear classifiers on input quantum mechanically encoded to improve feature extraction, which can be replaced by classical support vector machines if the encoding is classically tractable.
Such valid critical questions arise and should be addressed to fully comprehend and utilize the potential of quantum systems. Our survey aims to benefit researchers on these trails of thought as it is necessary to dive deeper into the field to come up with answers to such questions. Hence, our goal is to enable researchers using our survey to understand, comprehend, and compare the types of methodologies and tools they have to derive practical solutions and answers to their problems.
3.4 Preserving Maximum Information
@@
Quantum computers are machines that execute calculations using the characteristics of quantum mechanics. What is occurring at the quantum scale follows the different laws of physics from what is known empirically from classical physics. Due to the superposition phenomenon, a qubit can represent the dual state of 00 and $1$ simultaneously, while a classic binary bit can only take either 00 or $1$ at a time. Hence, to express n-bit combinations using a classical computer, we need $2^{n}$ combinations, which can be checked sequentially before the classical computer can find the solution. On the other hand, a quantum computer takes advantage of its wave nature so that the wave interference increases the probability of the desired state and decreases the other to reach a correct solution effectively all at once.
@@ 4.1 Categorization of QML Approaches
  • CQ refers to processing Classical data using Quantum machine learning algorithms and will be the main focus of this chapter.
@@ 4.2 Parametrized Quantum Circuits
A Variational Quantum Algorithm (VQA) combines a classical optimizer with a PQC. VQAs represent a promising approach to achieving quantum advantage over classical systems, even when using NISQ devices that come with various constraints. These constraints include a limited number of qubits, restricted qubit connectivity, and decoherence errors that limit the depth of quantum circuits (Cerezo et al., 2021). Despite these challenges, VQAs can leverage the strengths of both classical and quantum computing to tackle problems that are intractable for classical algorithms alone. Examples of applications in which VCAs excel are quantum chemistry, material science, financial modeling, and cryptography.
@@ 4.3 Quantum Optimization Problems
4.3.1 Variational Amplitude Amplification and Estimation
Variational Amplitude Amplification
Amplitude amplification is a technique in quantum computing algorithms derived by generalizing the working principle of Grover’s Search Algorithm (Grover, 1996). Initially discovered in Brassard & Hoyer , the approach to finding a solution using “amplitude amplification” gives rise to a new family of quantum algorithms with the advantage of obtaining a quadratic speedup over several classical algorithms in comparison to a brute-force search. However, to observe its best potential impact when applied to real-world problems that embody larger search spaces for solutions, we need heuristics along with "amplitude amplification" to make it practically useful (Brassard et al., 2002).
Variational Amplitude Estimation
Quantum Amplitude Estimation (QAE) is the task of finding an estimate for the amplitude of the solution state using the process of amplitude amplification described above. It can be generalized as a way of finding a solution to a combinatorial search problem, where components of interest in the space are projected in sub-spaces as a superposition of the required solution and other possibilities. Given the problem in this context, the process of “amplitude estimation” is the algorithm for estimating the probability that the final measurement after applying amplification will yield the desired state of the solution for the respective problem. The advantage of using this algorithm lies in the fact that it provides a quadratically faster solution than expected by a classical algorithm, provided the condition that the solution to the problem is known to be unique (Grover, 1996; 1997; Brassard et al., 2002). The first implementation of the algorithm in Brassard & Hoyer for amplitude estimation employs a strategy involving phase estimation, which makes the algorithm compute-intensive. Another variant employs an iterative strategy to avoid phase estimation, which represents a useful evolution of the amplitude estimation strategy (Grinko et al., 2021).
@@ 4.3.3 Variational Quantum Eigensolver
An instance of VQE requires the definition of two algorithmic sub-components, which are a quantum trial state (also called Ansatz (Zhang, 2012)) and a classical optimizer. The optimizer varies the parameters of the Anzatz, such that it works towards a state, as determined by its parameters, that results in the minimum expectation value being measured by the input operator (Hamiltonian). Additionally, the VQE can be used to model complex wave functions in polynomial time, making it one of the most promising NISQ applications for quantum computing. In practical applications, the VQE is used in chemistry for simulations of molecules, as well as logistics and network design.
@@ 4.3.4 Quantum Annealing
QA techniques require an objective function to be defined in the Ising model or Quadratic Unconstrained Binary Optimisation (QUBO). The Ising model is a mathematical model in statistical mechanics (Borle & Lomonaco, 2019). The QUBO (Jun, 2022) is mathematically equivalent to the Ising model and can be used to formulate a problem in a more simple way.
@@ 4.4.1 Quantum Convolutional Neural Networks
The structure of a classical CNN consists of applying alternating convolutional layers (with an activation function) and pooling layers, typically followed by fully-connected layers before the output is generated. In QCNNs, the convolution operations are performed as parameterized unitary rotations based on PQCs, executed on neighboring pairs of qubits. These convolutions are followed by pooling layers. The scope of pooling layers is to down-sample the results of the previous layer by reducing the data dimensions while inherently extracting the most relevant features. The dimensionality reduction with pooling layers reduces the dimensions of the quantum circuits, which in turn reduces the number of qubits in the circuit while retaining the maximum information possible from predecessor layers. Consequently, it reduces the computational cost of the complete circuit since the number of learnable parameters of the QCNN is equal to the qubits in the circuit. However, in quantum physics, it is not possible to directly remove qubits from the circuit. Hence, the pooling layers in quantum circuits are deployed by measuring a subset of the qubits and using these measurement results to control subsequent operations. The fully-connected layers are implemented as a multi-qubit operation on the remaining qubits before the final measurement. All the parameters involved in these operations are learned during training.
Figure 22: Functionality of a Quantum Convolutional Neural Network.
Figure 23: Functionality of a Quanvolutional Neural Network.
@@ 4.4.2 Quanvolutional Neural Networks
Considering the current limits of the quantum hardware, where the quantum circuit implementation of many qubits is challenging, QuanCNNs represent a valid solution for circumventing this obstacle since the deeper layers are still implemented in classical computing. The advantage of quantum computing is that it is used for implementing convolutional layers, where low-qubit circuits are sufficient to perform the required task well.
@@ 4.6 Quantum Kernels
Similarly to classical kernels, Quantum Kernels (QKs) help to deal with data points in high dimensional feature space and are typically employed to conduct unsupervised classification tasks. A quantum kernel can be defined as in Equation 3, where $x$ and $y$ are $n$-dimensional inputs and $Q(x,y)$ is the kernel function. $f$ denotes a function that performs the mapping from an $n$-dimensional space into an $m$-dimensional space, where $n<m$. $\langle f(x),f(y)\rangle$ denotes the dot product between the functions. This function serves as the foundation for performing clustering and regression algorithms in the quantum domain.
$Q(x,y)=\langle f(x),f(y)\rangle$ (3)
Constructing QKs for SVMs has been under investigation in recent works. The performance of the resulting quantum SVMs for classification problems with a small number of training points exceeds that of optimized classical models with conventional kernels (Torabian & Krems, 2023; Krunic et al., 2022). Moreover, QKs allow us to interpret the learning process with the help of quantum information tools (Sancho-Lorente et al., 2022).
Quantum clustering is a class of data clustering techniques that use the mathematical and conceptual tools of quantum mechanics. It belongs to a family of density-based clustering algorithms, where clusters are defined by areas of higher-density data points. Its design involves a scale-space probability function viewed at the lowest eigenstate of a Schrödinger equation, followed by simple analytic operations to derive a potential function whose minima determine cluster centers (see Figure 28). The design proposed by Horn & Gottlieb (2002) was applied to a two-dimensional system but is also scalable to higher dimensions.
@@
As demonstrated by Aïmeur et al. (2007), quantum clustering for different algorithms (divisive clustering, k-medians, and construction of a neighborhood graph) can obtain significant speedup compared to classical clustering.
@@ 4.8 Limitations of Current QML Approaches
  • The accuracy measured by QML algorithms is not yet matured to be on par with the state-of-the-art classical ML algorithms.
@@
  • The high-performance computation cost of developing large-scale QML incurs a high demand for access to tools and expensive resource infrastructures.
@@
  • For wide QNNs, the well-known Barren plateau problem (McClean et al., 2018) leads to trainability issues.
  • Since the parameter update of QML algorithms relies on classical optimization modules, the quantum technology cannot be utilized to its full extent.

### v4

@@ 3.1 Quantum Complexity Classes and Speedup
Quantum computing provides potential advantages over classical computing through its fundamentally different computational complexity characteristics. Classical complexity theory classifies problems into well-known classes such as P, NP, and PSPACE based on their solvability and computational resource requirements (Papadimitriou & Yannakakis, 1991) (see Appendix A for more details). Quantum computing expands this framework by introducing the complexity class Bounded-error Quantum Polynomial-time (BQP), consisting of problems efficiently solvable by quantum computers with bounded error probabilities (Watrous, 2008). As shown in Figure 13, the BQP class extends beyond classical P, potentially encompassing problems that are believed to be classically intractable. This perspective provides a formal basis for identifying computational problems where quantum algorithms may yield significant advantages.
3.2 Quantum Supremacy
“Quantum supremacy” refers to the demonstration that quantum computers can perform certain computational tasks significantly faster than any known classical algorithm (Frank Arute, 2019). Early demonstrations focused on random circuit sampling, generating bit-string outputs from a randomly chosen quantum circuit drawn from an ensemble believed to be hard to simulate classically (Boixo et al., 2018). The theoretical hardness of this task is grounded in complexity-theoretic conjectures (e.g., non-collapse of the polynomial hierarchy) rather than a proof that classical algorithms cannot ultimately match performance (Aaronson & Arkhipov, 2011). For instance, Google’s Sycamore quantum processor (Frank Arute, 2019) achieved a computational task in approximately $200$ seconds, whereas the estimated classical computation would require about $10,000$ years. Such demonstrations highlight the practical scenarios in which quantum devices exhibit exponential speedups over classical hardware, thus marking a critical milestone toward realizing practical quantum computing. However, subsequent advances in classical simulation, such as tensor network contraction optimizations and improved sampling algorithms, have dramatically reduced the estimated classical runtimes (Zhao et al., 2025). This evolution underscores that quantum supremacy benchmarks serve more as stress tests of classical simulation methods than as definitive proof of intractability.
3.3 Polynomial and Exponential Quantum Speedups
Quantum algorithms offer computational speedups over classical counterparts that are generally categorized as either exponential or polynomial. Exponential speedups, such as Shor’s algorithm for integer factorization (Shor, 1997), demonstrate drastic performance improvements, turning classically infeasible problems into quantum tractable tasks. In contrast, polynomial speedups, exemplified by Grover’s search algorithm (Grover, 1996), enhance performance significantly but do not reduce exponential complexity to polynomial. Understanding the types and implications of these speedups provides crucial insights for effectively leveraging quantum computing.
3.4 Practical Quantum Enhancement in Hybrid Quantum-Classical Approaches
In the NISQ era, quantum computing primarily complements classical computing rather than fully replacing it. Hybrid quantum-classical algorithms exploit quantum systems to accelerate specific sub-tasks within larger classical computations (Preskill, 2018). Recent work demonstrates quantum enhancements for practical computational problems, particularly in machine learning and optimization tasks (Pokharel & Lidar, 2023; Liu et al., 2021).
@@ 3.4 Practical Quantum Enhancement in Hybrid Quantum-Classical Approaches
In light of the hybrid quantum-classical systems, the work of Liu et al. (2021) discusses important aspects that should be changed to obtain meaningful developments and revolves around how heuristic-based algorithms do not provide formal evidence that showcases their genuine and consistent advantage over classical algorithms. The work of Kashif et al. (2025) demonstrates that HQNNs exhibit better scalability than classical neural networks when increasing classification task complexity. Moreover, the variational circuits can only implement linear classifiers on input quantum mechanically encoded to improve feature extraction, which can be replaced by classical support vector machines if the encoding is classically tractable. These hybrid strategies leverage quantum strengths in carefully selected areas, highlighting a realistic and immediate path toward practical quantum computing applications.
3.5 Quantum Information Encoding Efficiency
@@
Quantum computers leverage superposition and entanglement to represent and process complex datasets more compactly than classical systems. Due to the superposition phenomenon, a qubit can represent the dual state of 00 and $1$ simultaneously, while a classic binary bit can only take either 00 or $1$ at a time. Hence, to express n-bit combinations using a classical computer, we need $2^{n}$ combinations, which can be checked sequentially before the classical computer can find the solution. On the other hand, a quantum computer takes advantage of its wave nature so that the wave interference increases the probability of the desired state and decreases the other to reach a correct solution effectively all at once.
@@ 4.1 Categorization of QML Approaches
  • CQ refers to processing Classical data using Quantum machine learning algorithms and will be the main focus of this work.
@@ 4.2 Parametrized Quantum Circuits
A Variational Quantum Algorithm (VQA) combines a classical optimizer with a PQC. VQAs represent a promising approach to achieving quantum advantage over classical systems, even when using NISQ devices that come with various constraints. These constraints include a limited number of qubits, restricted qubit connectivity, and decoherence errors that limit the depth of quantum circuits (Cerezo et al., 2021a). Despite these challenges, VQAs can leverage the strengths of both classical and quantum computing to tackle problems that are intractable for classical algorithms alone. Examples of applications in which VCAs excel are quantum chemistry, material science, financial modeling, and cryptography.
@@ 4.3 Quantum Optimization Problems
4.3.1 Amplitude Amplification and Estimation
Amplitude amplification generalizes Grover’s search algorithm to boost the probability of measuring desired outcomes. This technique provides a quadratic speedup over classical brute-force search methods (Brassard & Hoyer, ; Grover, 1996).
Quantum Amplitude Estimation (QAE) builds upon amplitude amplification to estimate the probability amplitude of a target quantum state. The original version relies on phase estimation, which is resource-intensive, while iterative variants avoid phase estimation at the cost of more measurements (Brassard & Hoyer, ; Grinko et al., 2021). These methods are suited for estimating expectation values and integrals in quantum finance and chemistry.
@@ 4.3.3 Variational Quantum Eigensolver
An instance of VQE requires the definition of two algorithmic sub-components, which are a quantum trial state (also called Ansatz (Zhang, 2012)) and a classical optimizer. Common ansatz choices in VQE include hardware-efficient circuits and chemically inspired ones. A hardware-efficient ansatz uses layers of native quantum gates arranged in a repetitive pattern, enabling expressive states with relatively shallow circuits suited for NISQ devices (Kandala et al., 2017). In contrast, the Unitary Coupled Cluster (UCC) ansatz, popular in quantum chemistry, constructs the trial state by applying a series of particle excitation operators, producing a physically motivated but often deeper circuit (Romero et al., 2018). The optimizer varies the parameters of the Anzatz, such that it works towards a state, as determined by its parameters, that results in the minimum expectation value being measured by the input operator (Hamiltonian). Additionally, the VQE can be used to model complex wave functions in polynomial time, making it one of the most promising NISQ applications for quantum computing. In practical applications, the VQE is used in chemistry for simulations of molecules, as well as logistics and network design.
However, the performance of VQE in real-world settings is strongly influenced by the choice of ansatz and the optimization strategy. Deep or highly unstructured ansatze can lead to barren plateaus, i.e., vast flat regions in the parameter landscape where gradients vanish, hindering the training process (McClean et al., 2018; Cerezo et al., 2021b). The selection of a classical optimizer (e.g., gradient-based methods versus gradient-free heuristics) also impacts convergence. Certain optimizers may better navigate noise and avoid local minima, but no single optimizer works best for all problems (Tilly et al., 2022). Consequently, these factors have the following practical implications. An overly simplistic ansatz might not be expressive enough to reach the true ground state. On the other hand, an overly complex ansatz can require prohibitively deep circuits or become untrainable on noisy hardware, limiting VQE to relatively small systems in current experiments (Kandala et al., 2017; Tilly et al., 2022; Liu et al., 2024).
@@ 4.3.4 Quantum Annealing
QA techniques require an objective function to be defined in the Ising model or Quadratic Unconstrained Binary Optimisation (QUBO). The Ising model is a mathematical model in statistical mechanics (Borle & Lomonaco, 2018). The QUBO (Jun, 2022) is mathematically equivalent to the Ising model and can be used to formulate a problem in a more simple way.
@@
4.3.5 Key Distinctions
Unlike VQE, which uses gate-based quantum circuits and classical optimizers, QA uses adiabatic evolution and is often implemented on specialized hardware like D-Wave systems. While VQE targets general Hamiltonians using variational circuits and classical feedback loops, QA is a continuous-time analog process that exploits quantum tunneling and annealing to reach the solution. QA does not require gate-level programming, but lacks the circuit expressiveness and flexibility of VQE.
Both methods aim to approximate low-energy solutions, but differ in control mechanisms (discrete variational updates vs. adiabatic evolution) and hardware requirements. VQE is more tunable and programmable, while QA provides natural hardware-level implementation for combinatorial problems.
@@ 4.4.1 Quantum Convolutional Neural Networks
The structure of a classical CNN consists of applying alternating convolutional layers (with an activation function) and pooling layers, typically followed by fully-connected layers before the output is generated. In QCNNs, the convolution operations are performed as parameterized unitary rotations based on PQCs, executed on neighboring pairs of qubits. These convolutions are followed by pooling layers. The scope of pooling layers is to down-sample the results of the previous layer by reducing the data dimensions while inherently extracting the most relevant features. The dimensionality reduction with pooling layers reduces the dimensions of the quantum circuits, which in turn reduces the number of qubits in the circuit while retaining the maximum information possible from predecessor layers. Consequently, it reduces the computational cost of the complete circuit since the number of learnable parameters of the QCNN is equal to the qubits in the circuit. However, in quantum physics, it is not possible to directly remove qubits from the circuit. Hence, the pooling layers in quantum circuits are deployed by measuring a subset of the qubits and using these measurement results to control subsequent operations.
The fully-connected layers are implemented as a multi-qubit operation on the remaining qubits before the final measurement. All the parameters involved in these operations are learned during training.
@@ 4.4.2 Quanvolutional Neural Networks
Figure 22: Functionality of a Quantum Convolutional Neural Network.
Figure 23: Functionality of a Quanvolutional Neural Network.
QCNN vs. QuanCNNs: QCNNs implement an end-to-end deep network with a set of quantum circuits, whereas QuanCNNs use quantum circuits only for the convolutional feature extraction stage before feeding into classical layers. QCNNs leverage quantum pooling (via qubit measurement and feed-forward control) to progressively perform spatial down-sampling. However, a QCNN requires high qubit coherence for multiple sequential quantum layers, making large-depth QCNNs challenging on NISQ hardware. In contrast, QuanCNNs trade off full quantum depth for practicality. They insert one or a few quantum convolution layers as a preprocessing step to a classical CNN, exploiting quantum feature maps to encode richer representations while relying on classical layers for complex feature hierarchy. This hybrid design means QuanCNNs can work with shallow quantum circuits (low qubit counts and depths) that are feasible on current devices. In summary, QCNNs offer a more pure quantum architecture (potentially yielding quantum speedups or non-classical feature extraction) but face implementation limits due to noise and decoherence, whereas Quanvolutional CNNs are hybrid, easier to realize on near-term machines, and still enhance classical models by injecting quantum-computed features (often improving accuracy on image data with minimal quantum resources) (Zaman et al., 2024a).
@@ 4.6 Quantum Kernels
Quantum kernel methods extend this idea by leveraging quantum feature maps to encode data into a quantum Hilbert space (Havlíček et al., 2019). In a quantum kernel approach, each data point $\mathbf{x}$ is mapped to a normalized quantum state $|\phi(\mathbf{x})\rangle$ via a parameterized quantum circuit (feature map), and the kernel is defined as the overlap between two such states: $K(\mathbf{x},\mathbf{x}^{\prime})=|\langle\phi(\mathbf{x})|\phi(\mathbf{x}^{\prime})\rangle|$. By using quantum states as feature vectors, it is in principle possible to access extremely high-dimensional feature spaces and intricate entanglement correlations that have no efficient classical representation. A crucial consideration is to choose feature maps that are classically hard to simulate, as this is believed to be necessary for achieving quantum advantage in learning tasks (Havlíček et al., 2019). If the quantum kernel can be efficiently computed or approximated by a classical algorithm, then a classical model could replicate its results, negating any quantum advantage (Schuld & Killoran, 2019). Conversely, feature maps based on classically intractable quantum circuits (for example, instantaneous quantum polynomial (IQP) circuits) can yield kernel functions that separate data in ways unattainable by known classical kernels (Havlíček et al., 2019).
Designing these quantum feature maps comes with significant challenges. On one hand, highly expressive or deep quantum embeddings may suffer from phenomena analogous to barren plateaus, such as an exponential concentration of kernel values, where the inner products between states become almost uniformly small regardless of input, effectively making the kernel uninformative for learning (Schuld & Killoran, 2019). On the other hand, circuits complex enough to be classically non-simulable tend to be more sensitive to noise and harder to implement on near-term hardware. In practice, there is a trade-off between expressivity and practicality. The feature map must be complex enough to exploit uniquely quantum features of the data, yet structured enough to avoid excessive noise and to permit effective training. Ongoing research is exploring how to optimize quantum kernels under these constraints, including techniques to tailor or learn the feature map for a given problem (Torabian & Krems, 2023). This careful balance in quantum kernel design is crucial for their performance in real-world applications, where both quantum hardware limitations and the risk of classical simulability must be managed.
@@
Quantum clustering is a class of data clustering techniques that use the mathematical and conceptual tools of quantum mechanics. It belongs to a family of density-based clustering algorithms, where clusters are defined by areas of higher-density data points. Its design involves a scale-space probability function viewed at the lowest eigenstate of a Schrödinger equation, followed by simple analytic operations to derive a potential function whose minima determine cluster centers (see Figure 28). The design proposed by Horn & Gottlieb (2002) was applied to a two-dimensional system but is also scalable to higher dimensions. As demonstrated by Aïmeur et al. (2007), quantum clustering for different algorithms (divisive clustering, k-medians, and construction of a neighborhood graph) can obtain significant speedup compared to classical clustering.
@@ 4.8 Limitations of Current QML Approaches
  • The accuracy measured by QML algorithms is not yet mature enough to be on par with the state-of-the-art classical ML algorithms.
@@
  • The high computational cost of developing large-scale QML models incurs a high demand for specialized tools and expensive computational resources.
  • The performance of QML algorithms is fundamentally limited by current hardware capabilities (e.g., short qubit coherence times, limited qubit counts, and non-negligible error rates). Present processors offer only tens to a few hundred noisy qubits, with short coherence times and gate error rates often around $10^{-3}$ per operation, and many architectures have restricted qubit connectivity. These limitations severely cap the size and depth of quantum circuits that can be executed reliably. Significant advances in quantum hardware, such as longer coherence, lower error rates (via error mitigation or correction), and scaling up the number of qubits, are required for QML techniques to achieve their full potential.
@@
  • The barren plateau phenomenon, where gradients vanish exponentially with system size, significantly hinders the training of deeper or wider QML models such as VQEs, QCNNs, and quantum kernel methods (McClean et al., 2018). Mitigation strategies include problem-inspired ansatz design, local cost functions (Cerezo et al., 2021b), layer-wise training (Grant et al., 2019), entanglement management (Patti et al., 2021), residual connections (Kashif & Al-Kuwari, 2024), and adaptive initialization (Kashif et al., 2024; Zhuang & Guan, 2025).
  • Since the parameter update of QML algorithms relies on classical optimization modules, the quantum technology cannot be utilized to its full extent, because QML training is still bottlenecked by classical methods.
@@
  • Realizing a quantum advantage also demands that the data encoding (feature map) and ansatz used in QML are classically hard to simulate. If a QML model’s quantum states can be efficiently emulated by a classical algorithm, then the model cannot outperform classical methods (Thanasilp et al., 2024). In practice, variational circuits and quantum kernel methods must leverage feature maps that produce complex quantum states or entanglements which are infeasible to reproduce classically, as this is crucial for any genuine advantage over classical ML.
It is therefore evident that current QML techniques, such as VQEs, QNNs, and quantum kernel methods, will heavily rely on future improvements in quantum hardware. These approaches struggle with the shallow circuits and noise levels of today’s NISQ devices. Hence, better qubit coherence, higher qubit counts, and lower error rates will be essential to scale up QML algorithms. Ultimately, the viability of these algorithms in solving practical problems hinges on the advent of more robust quantum processors (potentially fault-tolerant devices) that can support deeper circuits and more qubits without being overwhelmed by errors. Concurrently, continued research is needed to design quantum data embeddings and model architectures that exploit quantum complexity beyond classical simulability, ensuring that QML models can capture patterns that classical ML cannot.